OpenStudy (valicat02):

The width of a rectangle is 8 cm less than its length. The perimeter of the rectangle is 52 cm. What is the width of the rectangle? A. 7 cm B. 9 cm C. 15 cm D. 17 cm

jhonyy9 (jhonyy9):

valicat02 how can you writing this given detail that the width is 8 cm less than the length of a rectangle ? using w to width and l to length

jhonyy9 (jhonyy9):

do you know what is formula of the perimeter of a rectangle ?

jhonyy9 (jhonyy9):

please collaborate - i like help you

jhonyy9 (jhonyy9):

how many sides has a rectangle ?

OpenStudy (valicat02):

4

jhonyy9 (jhonyy9):

yes sure 4 sides so from these 4 how many are length and how many are width ?

OpenStudy (valicat02):

2 length and 2 width?

jhonyy9 (jhonyy9):

yes exactly - very nice so than you know these details how you think please how will get the perimeter of the rectangle ?

OpenStudy (valicat02):

2w + 2l

jhonyy9 (jhonyy9):

very nice job - yes sure so the formula of perimeter of a rectangle is P = 2l +2w but you know that in this case the width is 8 cm less than the length of rectangle - what mean this - how can you writing this ?

OpenStudy (valicat02):

I'm not sure

jhonyy9 (jhonyy9):

so than you know that the length is l so 8 cm less than l how you can writing ? look please one example : 4 less than x will note x-4 ok. ?

jhonyy9 (jhonyy9):

using this example you need writing 8 less than l - so how will be ?

OpenStudy (valicat02):

l-8 ?

jhonyy9 (jhonyy9):

exactly so than you can writing that the width w = l - 8 yes ?

jhonyy9 (jhonyy9):

so than substitute this l-8 in place of w in the formula of perimeter how will get P = 2l +2w = ?

OpenStudy (valicat02):

P = 2l +2l-8

jhonyy9 (jhonyy9):

not exactly bc. there are P = 2l +2w = 2l +2(l-8) ok. bc. there is 2 multiply w what mean 2 multiply (l-8) right ?

OpenStudy (valicat02):

Yes

jhonyy9 (jhonyy9):

so can you distribute the parentheses ?

jhonyy9 (jhonyy9):

what mean 2time (l-8) so this mean 2 time l - 2 time 8 yes ?

jhonyy9 (jhonyy9):

so using this you can rewriting the formula of perimeter P = 2l +2(l-8) = 2l +2l -2*8 = ? continue please

jhonyy9 (jhonyy9):

2l +2l = ?

jhonyy9 (jhonyy9):

and 2 time 8 = ?

OpenStudy (valicat02):

2x8 is 16

jhonyy9 (jhonyy9):

yes sure and 2l +2l = ?

jhonyy9 (jhonyy9):

two apples plus two apples equal how many apples ?

OpenStudy (valicat02):

4

jhonyy9 (jhonyy9):

4 apples so than 2l plus 2l = how many l-es ?

jhonyy9 (jhonyy9):

than you assumed 2apples let being 2a plus 2a = 4a so than 2l +2l = ?

jhonyy9 (jhonyy9):

courage please this is easy

jhonyy9 (jhonyy9):

so than 2a +2a = 4a so 2l +2l = 4l right ?

jhonyy9 (jhonyy9):

in this case how you can continue this ? P = 2l +2(l-8) = 2l +2l -2*8 =

jhonyy9 (jhonyy9):

@TheSmartOne